# Worth Their Weight: Randomized and Regularized Block Kaczmarz Algorithms without Preprocessing

### v1

@@ Appendix C Proofs of ReBlocK Convergence Theorems
$\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{t}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{1}{2\lambda k\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$ (59)
Proof.
Since the ReBlocK iteration does not admit a simple orthogonal decomposition, we instead proceed by utilizing a bias-variance decomposition inspired by (Défossez & Bach, 2015; Jain et al., 2018b; Epperly et al., 2024).
@@
$\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq 2\mathbb{E}\,\left\lVert d_{t}\right\rVert^{2}+2\mathbb{E}\,\left\lVert v_{t}\right\rVert^{2}.$ (62)
@@
$\mathbb{E}\left[\left\lVert d_{s+1}^{2}\right\rVert|d_{s}\right]=d_{s}^{\top}\mathbb{E}\left[(I-P_{s})^{2}\right]d_{s}=d_{s}^{\top}(I-2\overline{P}+\mathbb{E}_{S\sim\rho}\left[P(S)^{2}\right])d_{s}.$ (63)
@@
$\mathbb{E}\left[\left\lVert d_{s+1}^{2}\right\rVert|d_{s}\right]\leq(1-\alpha)\mathbb{E}\,\left\lVert d_{s}\right\rVert^{2}.$ (65)
@@
$\mathbb{E}\,\left\lVert d_{t}\right\rVert^{2}\leq(1-\alpha)^{t}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}.$ (66)
@@
$\mathbb{E}\left[v_{s}\right]=(I-\overline{P})^{s}v_{0}=0.$ (67)
@@
$\mathbb{E}\left[\left\lVert v_{s+1}\right\rVert^{2}|v_{s}\right]$ $=v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))^{2}\right]v_{t}+2v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))A^{\top}W(S)r^{(\rho)}\right]+\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2}$
$\leq(1-\alpha)\left\lVert v_{s}\right\rVert^{2}+2v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))A^{\top}W(S)r^{(\rho)}\right]+\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2},$
@@
$\mathbb{E}\,\left\lVert v_{t}\right\rVert^{2}\leq\frac{1}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2}.$ (68)
@@
$\left\lVert A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}\right\rVert\leq\frac{1}{2\sqrt{\lambda k}}.$ (70)
@@
$\mathbb{E}\,\left\lVert v_{t}\right\rVert^{2}\leq\frac{1}{4\lambda k\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$ (71)
Combining the bias bound (66) and the variance bound (71) using (62) yields the desired result. ∎
Proof of Theorem 4.1.
The result will follow from the appropriate application of Lemmas A.1, A.4 and C.1. To apply these lemmas it is first required to verify y $\overline{P}\preceq I$ and $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$. For the first result, it suffices to observe that $P(S)\preceq P_{RBK}(S)\preceq I$ for all $S$.
@@
$x^{\top}\overline{P}x=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[x^{\top}A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}A_{S}x\right]\geq\frac{k}{m}\cdot\frac{(a_{i}^{\top}x)^{2}}{\left\lVert AA^{\top}+k\lambda I\right\rVert_{2}}>0,$ (72)
@@
$\mathbb{E}\left[x_{T}\right]-x^{(\rho)}=(I-\overline{P})^{T}(x_{0}-x^{(\rho)}).$ (73)
@@
$\left\lVert\mathbb{E}\left[x_{T}\right]-x^{(\rho)}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{(\rho)}\right\rVert.$ (74)
using Lemmas A.2 and A.3 and the proven properties of $\overline{P}.$
@@
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x^{(\rho)}\right\rVert^{2}+\frac{1}{2\lambda k\alpha^{2}(T-T_{b})}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$ (75)
@@
$\left\lVert A_{S}A_{S}^{\top}\right\rVert=\left\lVert A_{S}^{\top}A_{S}\right\rVert=\left\lVert\sum_{i\in S}a_{i}a_{i}^{\top}\right\rVert\leq\sum_{i\in S}\left\lVert a_{i}a_{i}^{\top}\right\rVert\leq Nk.$ (76)
@@
$\lambda kI\preceq A_{S}A_{S}^{\top}+\lambda kI\preceq(N+\lambda)kI.$ (77)
@@
$\overline{W}=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[I_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}I_{S}\right]\succeq\frac{1}{(N+\lambda)m}I,$ (78)
$\overline{W}=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[I_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}I_{S}\right]\preceq\frac{1}{\lambda m}I.$ (79)
@@
$\kappa(\overline{W})\leq 1+\frac{N}{\lambda},$ (80)
and the final claim of the theorem follows by Lemma A.5. ∎
@@ Appendix D Noisy Linear Least-squares
For such problems with increasing numbers of rows, both RBK and ReBlocK converge (in a Monte Carlo sense) to the ordinary least-squares solution rather than a weighted solution. This can be viewed as an advancement over previous results on noisy linear systems, such as (Needell, 2010), which show only convergence to a finite horizon even for arbitrarily large noisy systems. The advance comes from treating the noise model explicitly and applying tail averaging to the algorithm.
@@
$\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$ (82)
@@
$\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$ (83)
Unified proof of Theorems D.1 and D.2.
@@
$y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{a\sim\mathbf{A},z\sim\mathbf{Z}}\left[(a^{\top}(x-y)-z)^{2}\right].$ (84)
Furthermore, the ordinary least-squares solution $x^{*}$ is the corresponding empirical risk minimizer. It is a standard result in asymptotic statistics that as the number of samples $m$ grows, the empirical minimizer $x^{*}$ converges to the true solution $y$ (Lehmann & Casella, 2006; Van der Vaart, 2000).
@@
$\overline{P}x^{(\rho)}=A^{\top}\overline{W}b=\overline{P}x^{*}+A^{\top}\overline{W}r.$ (85)
@@
$\lim_{m\rightarrow\infty}\overline{P}^{+}A^{\top}\overline{W}r=0.$ (87)
@@
$\lim_{m\rightarrow\infty}\overline{P}=\lim_{m\rightarrow\infty}\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[P(A_{S})\right]=\mathbb{E}_{A_{S}\sim\mathbf{A}^{k}}\left[P(A_{S})\right]\coloneqq\hat{P},$ (88)
$\lim_{m\rightarrow\infty}\overline{P}^{+}A^{\top}\overline{W}r=\hat{P}^{+}\lim_{m\rightarrow\infty}\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[A_{S}^{\top}M(A_{S})r_{S}\right]=\hat{P}^{+}\mathbb{E}_{A_{S}\sim\mathbf{A}^{k}}\left[A_{S}^{\top}M(A_{S})\right]\mathbb{E}_{r_{S}\sim\mathbf{Z}^{k}}\left[r_{S}\right]=0.$ (89)
@@
$\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$ (90)
@@ Appendix E Sampling from a Determinantal Point Process
In Section 4, we have presented the convergence of ReBlocK to a weighted least-squares solution under the uniform sampling distribution $\rho=\mathbf{U}(m,k)$. However, the explicit convergence rate $1-\alpha$ is still unknown. Motivated by recent work on Kaczmarz algorithms with determinantal point process (DPP) sampling (Dereziński & Yang, 2024), we now show that under DPP sampling, ReBlocK converges to the ordinary least-squares solution and the convergence of ReBlocK can have a much better dependence on the singular values of $A$ than mSGD.
@@
$\Pr[S]=\frac{\det(A_{S}A_{S}^{\top}+\lambda kI)}{\sum_{|S^{\prime}|=k}\det(A_{S^{\prime}}A_{S^{\prime}}^{\top}+\lambda kI)}.$ (91)
Our analysis is similar to parts of the nearly concurrent work (Dereziński et al., 2025), though (Dereziński et al., 2025) does not incorporate a fixed size $k$ for their regularized DPP distribution..
@@
$\left\lVert\mathbb{E}\left[x_{T}\right]-x^{*}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{*}\right\rVert.$ (92)
@@
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{*}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x_{0}-x^{*}\right\rVert^{2}\\ +\frac{1}{2\lambda\alpha^{2}(T-T_{b})}\cdot\max_{i}(r_{i}^{2}).$ (93)
@@
$\alpha\geq\frac{k-\ell}{(k-\ell)+\kappa_{\ell}^{2}(A)+(m+k-2\ell)\frac{\lambda}{\sigma_{n_{r}}^{2}}},$ (94)
@@
$\begin{split}\alpha&\geq\frac{k}{k+2\kappa_{k/2}^{2}(A)+2m\lambda/\sigma_{n}^{2}}\\ &\approx\frac{k}{2\kappa^{2}_{k/2}(A)+2\lambda\kappa_{\rm dem}^{2}(A)}.\end{split}$ (95)
@@
Our bound for the variance of the tail-averaged estimator is crude and could likely be improved using techniques such as those of (Dereziński et al., 2025). Nonetheless, it is interesting to note that the dependence of the variance on the residual vector can be worse in the case of DPP sampling than in the case of uniform samping. This is because the DPP distribution can potentially sample the residual vector in unfavorable ways. It is thus unclear whether DPP sampling is a good strategy for arbitrary inconsistent systems, even if it can be implemented efficiently
Proof.
@@
We begin by verifying that $x^{(\rho)}=x^{*}$. For a vector $u\in\mathbb{R}^{m}$, we denote its elementary symmetric polynomial by $p_{k}(u):=\sum_{S\in\binom{[m]}{k}}\prod_{i\in S}u_{i}$, where $\binom{[m]}{k}$ is the set of all subsets of size $k$ from the set $[m]=\{1,\ldots,m\}$. Applying equation (5.3) of (Dereziński & Yang, 2024) by using $B=AA^{\top}+\lambda I$ leads to
$\overline{W}=\mathbb{E}_{S\sim\rho}\left[I_{S}^{\top}(I_{S}(AA^{\top}+\lambda I)I_{S}^{\top})^{-1}I_{S}\right]=\frac{U{\rm diag}(p_{k-1}(q_{-1}),\ldots,p_{k-1}(q_{-m}))U^{\top}}{p_{k}(q)},$ (96)
@@
$\overline{P}=A^{\top}\overline{W}A=V\Sigma^{\top}\frac{{\rm diag}(p_{k-1}(q_{-1}),\ldots,p_{k-1}(q_{-m}))}{p_{k}(q)}\Sigma V^{\top}.$ (97)
We now follow closely the logic of the proof of Lemma 4.1 of (Dereziński & Yang, 2024), making appropriate modifications to handle the fact that in GRK we invert $AA^{\top}+\lambda I$ rather than just $AA^{\top}$. Note also that through (97) we have explicitly verified that $\operatorname{range}(\overline{P})=\operatorname{range}(A^{T})$, which is the needed condition for the contraction properties to hold in our convergence bound.
@@
$B_{\ell}=U{\rm diag}\left(q_{1},\ldots q_{m}\right)U^{\top}+\frac{1}{k-\ell}\sum_{j>\ell}^{m}q_{j}I.$ (98)
Note that we have used a mildly simpler and looser approximation than (Dereziński & Yang, 2024) by replacing $\frac{k-\ell-1}{k-\ell}$ with $1$ in the first term.
The same arguments as in (Dereziński & Yang, 2024) then imply that
$\alpha=\sigma_{\min}^{+}(\overline{P})\geq\sigma_{\min}^{+}(A^{\top}B_{\ell}^{-1}A).$ (99)
@@
$\frac{1}{2\lambda k\alpha^{2}(T-T_{b})}\mathbb{E}\,_{S\sim\rho}\left\lVert r_{S}\right\rVert^{2}$ (100)
which is esentially the same as for ReBlocK-U but with the ordinary least-squares residual $r$ instead of the weighted residual $r^{(\rho)}$. Unfortunately, in the case of DPP sampling, we have no guarantee on how $S$ samples the residual vector $r^{(\rho)}$. Thus, the best we can do is uniformly bound $\left\lVert r_{S}\right\rVert^{2}\leq k\cdot\max_{i}(r_{i}^{2})$. This yields the bound in the theorem. ∎

### v2

@@ Appendix C Proofs of ReBlocK Convergence Theorems
(60) $\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{t}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{1}{2\lambda k\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$
Proof C.2.
Since the ReBlocK iteration does not admit a simple orthogonal decomposition, we instead proceed by utilizing a bias-variance decomposition inspired by [12, 23, 18].
@@
(63) $\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq 2\mathbb{E}\,\left\lVert d_{t}\right\rVert^{2}+2\mathbb{E}\,\left\lVert v_{t}\right\rVert^{2}.$
@@
(64) $\mathbb{E}\left[\left\lVert d_{s+1}^{2}\right\rVert|d_{s}\right]=d_{s}^{\top}\mathbb{E}\left[(I-P_{s})^{2}\right]d_{s}=d_{s}^{\top}(I-2\overline{P}+\mathbb{E}_{S\sim\rho}\left[P(S)^{2}\right])d_{s}.$
@@
(66) $\mathbb{E}\left[\left\lVert d_{s+1}^{2}\right\rVert|d_{s}\right]\leq(1-\alpha)\mathbb{E}\,\left\lVert d_{s}\right\rVert^{2}.$
@@
(67) $\mathbb{E}\,\left\lVert d_{t}\right\rVert^{2}\leq(1-\alpha)^{t}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}.$
@@
(68) $\mathbb{E}\left[v_{s}\right]=(I-\overline{P})^{s}v_{0}=0.$
@@
$\mathbb{E}\left[\left\lVert v_{s+1}\right\rVert^{2}|v_{s}\right]=$ $v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))^{2}\right]v_{t}+2v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))A^{\top}W(S)r^{(\rho)}\right]$
$+\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2}$
$\leq$ $(1-\alpha)\left\lVert v_{s}\right\rVert^{2}+2v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))A^{\top}W(S)r^{(\rho)}\right]$
$+\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2},$
@@
(69) $\mathbb{E}\,\left\lVert v_{t}\right\rVert^{2}\leq\frac{1}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2}.$
@@
(71) $\left\lVert A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}\right\rVert\leq\frac{1}{2\sqrt{\lambda k}}.$
@@
(72) $\mathbb{E}\,\left\lVert v_{t}\right\rVert^{2}\leq\frac{1}{4\lambda k\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$
Combining the bias bound (67) and the variance bound (72) using (63) yields the desired result.
Proof C.3 (Proof of Theorem 4.1).
The result will follow from the appropriate application of Lemmas A.1, A.7, and C.1. To apply these lemmas it is first required to verify y $\overline{P}\preceq I$ and $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$. For the first result, it suffices to observe that $P(S)\preceq P_{RBK}(S)\preceq I$ for all $S$.
@@
(73) $x^{\top}\overline{P}x=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[x^{\top}A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}A_{S}x\right]\geq\frac{k}{m}\cdot\frac{(a_{i}^{\top}x)^{2}}{\left\lVert AA^{\top}+k\lambda I\right\rVert_{2}}>0,$
@@
(74) $\mathbb{E}\left[x_{T}\right]-x^{(\rho)}=(I-\overline{P})^{T}(x_{0}-x^{(\rho)}).$
@@
(75) $\left\lVert\mathbb{E}\left[x_{T}\right]-x^{(\rho)}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{(\rho)}\right\rVert.$
using Lemmas A.3 and A.5 and the proven properties of $\overline{P}.$
@@
(76) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x^{(\rho)}\right\rVert^{2}+\frac{1}{2\lambda k\alpha^{2}(T-T_{b})}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$
@@
(77) $\left\lVert A_{S}A_{S}^{\top}\right\rVert=\left\lVert A_{S}^{\top}A_{S}\right\rVert=\left\lVert\sum_{i\in S}a_{i}a_{i}^{\top}\right\rVert\leq\sum_{i\in S}\left\lVert a_{i}a_{i}^{\top}\right\rVert\leq Nk.$
@@
(78) $\lambda kI\preceq A_{S}A_{S}^{\top}+\lambda kI\preceq(N+\lambda)kI.$
@@
(79) $\overline{W}=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[I_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}I_{S}\right]\succeq\frac{1}{(N+\lambda)m}I,$
(80) $\overline{W}=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[I_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}I_{S}\right]\preceq\frac{1}{\lambda m}I.$
@@
(81) $\kappa(\overline{W})\leq 1+\frac{N}{\lambda},$
and the final claim of the theorem follows by Lemma A.9.
@@ Appendix D Noisy Linear Least-squares
For such problems with increasing numbers of rows, both RBK and ReBlocK converge (in a Monte Carlo sense) to the ordinary least-squares solution rather than a weighted solution. This can be viewed as an advancement over previous results on noisy linear systems, such as [30], which show only convergence to a finite horizon even for arbitrarily large noisy systems. The advance comes from treating the noise model explicitly and applying tail averaging to the algorithm.
@@
(83) $\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$
@@
(84) $\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$
Proof D.3 (Unified proof of Theorems D.1 and D.2).
@@
(85) $y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{a\sim\mathbf{A},z\sim\mathbf{Z}}\left[(a^{\top}(x-y)-z)^{2}\right].$
Furthermore, the ordinary least-squares solution $x^{*}$ is the corresponding empirical risk minimizer. It is a standard result in asymptotic statistics that as the number of samples $m$ grows, the empirical minimizer $x^{*}$ converges to the true solution $y$ [25, 44].
@@
(86) $\overline{P}x^{(\rho)}=A^{\top}\overline{W}b=\overline{P}x^{*}+A^{\top}\overline{W}r.$
@@
(88) $\lim_{m\rightarrow\infty}\overline{P}^{+}A^{\top}\overline{W}r=0.$
@@
(89) $\lim_{m\rightarrow\infty}\overline{P}=\lim_{m\rightarrow\infty}\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[P(A_{S})\right]=\mathbb{E}_{A_{S}\sim\mathbf{A}^{k}}\left[P(A_{S})\right]\coloneqq\hat{P},$
$\lim_{m\rightarrow\infty}\overline{P}^{+}A^{\top}\overline{W}r$ $=\hat{P}^{+}\lim_{m\rightarrow\infty}\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[A_{S}^{\top}M(A_{S})r_{S}\right]$
$=\hat{P}^{+}\mathbb{E}_{A_{S}\sim\mathbf{A}^{k}}\left[A_{S}^{\top}M(A_{S})\right]\mathbb{E}_{r_{S}\sim\mathbf{Z}^{k}}\left[r_{S}\right]=0.$
@@
(90) $\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$
@@ Appendix E Sampling from a Determinantal Point Process
In Section 4, we have presented the convergence of ReBlocK to a weighted least-squares solution under the uniform sampling distribution $\rho=\mathbf{U}(m,k)$. However, the explicit convergence rate $1-\alpha$ is still unknown. Motivated by recent work on Kaczmarz algorithms with determinantal point process (DPP) sampling [15], we now show that under DPP sampling, ReBlocK converges to the ordinary least-squares solution and the convergence of ReBlocK can have a much better dependence on the singular values of $A$ than mSGD.
@@
(91) $\Pr[S]=\frac{\det(A_{S}A_{S}^{\top}+\lambda kI)}{\sum_{|S^{\prime}|=k}\det(A_{S^{\prime}}A_{S^{\prime}}^{\top}+\lambda kI)}.$
Our analysis is similar to parts of the nearly concurrent work [16], though [16] does not incorporate a fixed size $k$ for their regularized DPP distribution..
@@
(92) $\left\lVert\mathbb{E}\left[x_{T}\right]-x^{*}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{*}\right\rVert.$
@@
(93) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{*}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x_{0}-x^{*}\right\rVert^{2}\\ +\frac{1}{2\lambda\alpha^{2}(T-T_{b})}\cdot\max_{i}(r_{i}^{2}).$
@@
(94) $\alpha\geq\frac{k-\ell}{(k-\ell)+\kappa_{\ell}^{2}(A)+(m+k-2\ell)\frac{\lambda}{\sigma_{n_{r}}^{2}}},$
@@
(95) $\begin{split}\alpha&\geq\frac{k}{k+2\kappa_{k/2}^{2}(A)+2m\lambda/\sigma_{n}^{2}}\\ &\approx\frac{k}{2\kappa^{2}_{k/2}(A)+2\lambda\kappa_{\rm dem}^{2}(A)}.\end{split}$
@@
Our bound for the variance of the tail-averaged estimator is crude and could likely be improved using techniques such as those of [16]. Nonetheless, it is interesting to note that the dependence of the variance on the residual vector can be worse in the case of DPP sampling than in the case of uniform samping. This is because the DPP distribution can potentially sample the residual vector in unfavorable ways. It is thus unclear whether DPP sampling is a good strategy for arbitrary inconsistent systems, even if it can be implemented efficiently
Proof E.2.
@@
We begin by verifying that $x^{(\rho)}=x^{*}$. For a vector $u\in\mathbb{R}^{m}$, we denote its elementary symmetric polynomial by $p_{k}(u):=\sum_{S\in\binom{[m]}{k}}\prod_{i\in S}u_{i}$, where $\binom{[m]}{k}$ is the set of all subsets of size $k$ from the set $[m]=\{1,\ldots,m\}$. Applying equation (5.3) of [15] by using $B=AA^{\top}+\lambda I$ leads to
(96) $\overline{W}=\mathbb{E}_{S\sim\rho}\left[I_{S}^{\top}(I_{S}(AA^{\top}+\lambda I)I_{S}^{\top})^{-1}I_{S}\right]=\frac{U{\rm diag}(p_{k-1}(q_{-1}),\ldots,p_{k-1}(q_{-m}))U^{\top}}{p_{k}(q)},$
@@
(97) $\overline{P}=A^{\top}\overline{W}A=V\Sigma^{\top}\frac{{\rm diag}(p_{k-1}(q_{-1}),\ldots,p_{k-1}(q_{-m}))}{p_{k}(q)}\Sigma V^{\top}.$
We now follow closely the logic of the proof of Lemma 4.1 of [15], making appropriate modifications to handle the fact that in GRK we invert $AA^{\top}+\lambda I$ rather than just $AA^{\top}$. Note also that through (97) we have explicitly verified that $\operatorname{range}(\overline{P})=\operatorname{range}(A^{T})$, which is the needed condition for the contraction properties to hold in our convergence bound.
@@
(98) $B_{\ell}=U{\rm diag}\left(q_{1},\ldots q_{m}\right)U^{\top}+\frac{1}{k-\ell}\sum_{j>\ell}^{m}q_{j}I.$
Note that we have used a mildly simpler and looser approximation than [15] by replacing $\frac{k-\ell-1}{k-\ell}$ with $1$ in the first term.
The same arguments as in [15] then imply that
(99) $\alpha=\sigma_{\min}^{+}(\overline{P})\geq\sigma_{\min}^{+}(A^{\top}B_{\ell}^{-1}A).$
@@
(100) $\frac{1}{2\lambda k\alpha^{2}(T-T_{b})}\mathbb{E}\,_{S\sim\rho}\left\lVert r_{S}\right\rVert^{2}$
which is esentially the same as for ReBlocK-U but with the ordinary least-squares residual $r$ instead of the weighted residual $r^{(\rho)}$. Unfortunately, in the case of DPP sampling, we have no guarantee on how $S$ samples the residual vector $r^{(\rho)}$. Thus, the best we can do is uniformly bound $\left\lVert r_{S}\right\rVert^{2}\leq k\cdot\max_{i}(r_{i}^{2})$. This yields the bound in the theorem.